Circular No. 7744

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001fg IN ANONYMOUS GALAXY D. Vanden Berk and B. Wilhite, Experimental Astrophysics Group,

Fermi National Accelerator Laboratory, report the detection of an apparent supernova by the Sloan Digital Sky Survey (SDSS) in images taken with the SDSS 2.5-m telescope on Oct. 15.15 UT. Magnitudes derived for the new object are u = 21.16, g = 19.20, r = 17.87, i = 18.17, z = 17.86 (dereddened for Galactic extinction using the values given by Schlegel et al. 1998, Ap.J. 500, 525). SN 2001fg is located at α = 21^{h}12^{m}45^{s}.23, δ = −0^{o}52′36′′.2 (equinox 2000.0), which is 3′′ west and 4′′ south of the nucleus of a galaxy having position end figures 45^{s}.45, 32′′.3. Follow-up observations taken at the Apache Point Observatory 3.5-m telescope on Oct. 31.1 by A. Miceli, C. Stubbs, and B. Lawton confirm the existence of the new star. SN 2001fg is not present in images taken by the SDSS on June 15.3 (limiting mag r ≈ 23.0).

SUPERNOVA 2001fh IN PGC 66592 P. Martin and W. D. Li, University of California at Berkeley, report

the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Nov. 3.2 (mag ∼ 17.2) and 5.2 UT (mag ∼ 16.9). SN 2001fh is located at α = 21^{h}20^{m}42^{s}.50, δ = +44^{o}23′53′′.2 (equinox 2000.0), which is 0′′.6 east and 6′′.3 south of the nucleus of PGC 66592. A KAIT image taken on Oct. 27.2 showed nothing at this position (limiting mag ∼ 19.5).

NOVA IN THE SMALL MAGELLANIC CLOUD 2001 W. Liller, Viña del Mar, Chile, provides the following improved posi-

tion for this nova (IAUC 7738): α = 0^{h}46^{m}27^{s}.56 ± 0^{s}.19, δ = −73^{o}29′46′′.2 ± 0′′.6 (equinox 2000.0).

G. L. Bosch, R. H. Barbá, and N. I. Morrell, Observatorio Astronómico, Universidad Nacional de La Plata, report that échelle spectroscopy (range 460–780 nm, resolution 0.04 nm) obtained of the nova at CASLEO Observatory on Oct. 26.149 UT shows strong, low-excitation emission lines of H, Fe II, [O I], O I, Na I, and possibly [Ni I] with broad P-Cyg profiles. Terminal wind velocities measured from the absorption-profile edge in Hβ indicates expansion of up to 1800 km/s, while Fe II and Na I lines show a slower expansion velocity of 950 km/s. The FWHM of the Hα emission was ∼ 1400 km/s.

2001 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green